OpenStudy (anonymous):

I WILL GIVE YOU A MEDAL ANYONE Evaluate Dx. -4x + 7y = 1 25 = 2x + 5y

OpenStudy (gorv):

wat is Dx???

OpenStudy (anonymous):

He probally means just x..

OpenStudy (gorv):

-4x + 7y = 1 ----------------- 1 2x + 5y=25 --------------------- 2 multiply 2 by 2 on both side and add in equ 1 4x+10y=50 -4x+7y=1 on adding 10y+7y=51 17y=51 y=3 x=5

OpenStudy (gorv):

@shortie155

OpenStudy (anonymous):

@gorv